1. A two-way ANOVA was performed to assess the effects of species and seed density on the growth of seedlings from the Brassica family. See JMP output Analysis of Variance Sum of 5 29134312 6 1403.2611 C. Total51 4316.6923 SourceDF Squares Mean Square Ratic Model Error 582.686 19.1009 0.506 Prob> F 0001 Effect Tests Sum of Source Seed density Species Species Seed density Nparm DF Squares FRatio Prob> F 0001 1 10.0323 0.0011 0.9742 2 211.2774 0.1848 0.8319 2 2 29013800 a. How many seed density levels were there? b. How many species of seedlings were used? How many observations were in the experiment, total? c. Calculate the F statistic for seed density. Hint: you will need to calculate a mean square first. MS -SS/df
Answer #1

a)
number of seed density = df +1 = 2+1 = 3
b)
number of species = 1+1=2
total observation =51+1=52

c)
for seed density
MS= SS/df=2901.38/2=1450.69
F = MS / MS_error = 1450.69/30.506 = 47.5542516

Please rate